OpenStudy (lgg23):

Evaluate the limit as x approaches positive infinity: sqrt(X^2 - 6x +1)-x

OpenStudy (lgg23):

\[\lim_{x \rightarrow \infty} \sqrt{x^{2}-6x+1}-x\]

OpenStudy (anonymous):

answer is -3 and i can show at least the first step

OpenStudy (anonymous):

multiply top and bottom as if you were rationalizing the denominator by \[\sqrt{x^2-6x+1}+x\] \[\frac{x^2-6x+1-x^2}{\sqrt{x^2-6x+1}+x}\] combine like terms in the numerator to get \[\frac{-6x+1}{\sqrt{x^2-6x+1}+x}\] and now let x go to infinity. you will get \[\frac{-6}{2}=-3\] but the details for that step are up to you. basically the only terms that matter in the denominator are \[\sqrt{x^2}\] and \[x\]

OpenStudy (anonymous):

i have seen this written out where you divide each term by x. that may be a good explanation, with the sophistication that if you divide by x inside the radical you you have to write it as \[\frac{1}{x^2}\]so the denominator will look something like \[\sqrt{\frac{x^2}{x^2}-\frac{6x}{x^2}+\frac{1}{x^2}}+\frac{x}{x}\] or \[\sqrt{1-\frac{6}{x}+\frac{1}{x^2}}+1\] and the numerator will look like \[-\frac{6x}{x}+\frac{1}{x}=-6+\frac{1}{x}\]

OpenStudy (anonymous):

then let x go to infinity, argue that each term with an "x" in the denominator goes to zero, and you are left with \[-\frac{6}{2}\]